2. If assume collection of modules have exponentially distributed lifetimes (age of component doesn't matter in failure probability) and modules fail independently, overall failure rate of collection is sum of failure rates of modules. Calculate MTTF of a disk subsystem with • 10 disks, each rated at 1,000,000 hour MTTF • 1 SCSI controller, 500,000 hour MTTF • 1 power supply, 200,000 hour MTTF • 1 fan, 200,000 MTTF • 1 SCSI cable, 1,000,000 hour MTTF 1) Failure Rate? 2) MTTF? 3) What if 1000,000 MTTF > 100 years, Make an example for this condition.
Answer #1

Answer: ! + failure rate Tox 1,000,000 500,000 1 ( t t + 200,000 2007000 1,000,000 23 = C1042+5+571) 1,000,000 ,000,000 MTTF- MTTF (Human) 100,000 =5000.XY3 20 Deathe >85years olds - 20,000/100,000 MTTF Chumain) ) = 100.000 /20.000 = 5 years 1,000,Hardware M7 rf is often 100,000 to 1,000,000 hrs 8 MITF is often I tolo * hours there is good chances that it one component f